OpenStudy (anonymous):

lim as x approaches -2 from the left side. 2x^2+x+1 / x+2

OpenStudy (nikvist):

\[\lim_{x\rightarrow -2-}\frac{2x^2+x+1}{x+2}=\frac{7}{0-}=-\infty\]